what is the H+ in a solution with a pH of 9.28


1
Expert's answer
2021-05-25T07:48:31-0400

pH=-log10[H+(aq)]

-pH=log10[H+(aq)]

[H+(aq)]=10-pH

[H+(aq)]=10-9.28

[H+(aq)]=5.25*10-10 moldm-3